OpenStudy (anonymous):

which of the following is false? A.5+7=7+5 B.-13/37.square root of 43=square root of 43 (-13/37) C.3x^2+14y=14y+3x^2 D.6x^3-9x=9x-6x^3

OpenStudy (jwheele1):

.5+7=7+5 12 = 12 true

OpenStudy (jwheele1):

so not a

OpenStudy (jwheele1):

6x^3-9x=9x-6x^3 make x a number.....lets say 0. 6(0)-9(0)=9(0)-6(0) 0 -0 = 0 - 0 0 = 0 true so not D

OpenStudy (jwheele1):

3x^2+14y=14y+3x^2 make x a number (0) make y a different number (1) 3(0)+14(1)=14(1)+3(0) 0 + 14 = 14 +0 14 = 14 true so not C

OpenStudy (jwheele1):

rephrase b please

OpenStudy (anonymous):

-13/37 times\[\sqrt{43}\=sqrt43(-13/37)

OpenStudy (jwheele1):

got it

OpenStudy (jwheele1):

that one is also true because of the commutative property of multiplication....

OpenStudy (jwheele1):

somethign is just not right here...lol

Directrix (directrix):

@jwheele1 OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (jwheele1):

@Directrix apologize, I thought that's what I was doing by working out the steps. Ill wait for a reply for each step in the future. I do not condone giving a straight forward answer ever.

Directrix (directrix):

@jwheele1 I hear ya. It is tough and frustrating to help the Askers when they post a question and disappear. When the Asker is "live" on the thread, some are shy and will not respond.